From: STOSC::SMM "On-Duty SMM x5074" 29-JAN-1993 19:29:57.70 To: @[SMM.DIS]HST_STAT CC: Subj: HST STATUS REPORT, 01/28 - 01/29 HUBBLE SPACE TELESCOPE DAILY REPORT #812 PERIOD COVERED: 0800 EST 01/28/93 - 0800 EST 01/29/93 Daily Status Report as of 029/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed FOS/RD 3296 (Supernova Remnants and Nucleo Synthesis) The Faint Object Spectrograph was used on both the red and blue sides to make Series of spectral observations of the supernova remnant NGC-4449-SN. The binary target acquisition failed on the first set, and a status buffer error message indicated no target. Although the subsequent peakup was "successful", the science data for the peakup showed no target. An HSTAR was filed. In the next series of eight observations the spectra showed broad emission lines. The last three observations showed no real target, and the spectra showed only noise. 1.2 Completed FOS/RD 4206 (Spectroscopy of Super Metal Poor Galaxies) The Faint Object Spectrograph was used to make three 2400 second exposures of the super metal poor irregular galaxy GR-8. The three observations were received and displayed. The first binary target acquisition failed, and generated a status buffer error message. A corrective SIP maneuver was uplinked since the target acquisition failed. The science observations were reported to mostly noise. An HSTAR was written. 1.3 Completed WFC/PC 4207 (Photometric Calibration Monitor) Both the Wide Field Camera and the Planetary Camera were used with six wide band filters to obtain single chip readouts of W2 and P6 to measure the ultraviolet flux of standard star BD+75D325. All six images were received and displayed. They all were reported to look good. 1.4 Completed FOS/BL 4698 (FOS Dark Monitoring) The Faint Object Spectrograph was used on both the red and blue sides to obtain dark exposures to determine the instrumental background. The observations were executed as scheduled, and no problems were reported. 1.5 Completed PC 4140 (WF/PC Calibration:Rapid Internal Monitor - Part 2) The Planetary Camera was used to obtain internal flat observations to check for new forms of contamination. The OSS monitored to observations, and no problems were reported. 1.6 Uplinked Updated Ephemeris Tables, Per MOSES 2871 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 2 Scheduled Reacquisitions : 9 Successful : 9 Losses of Lock : 0 2.2 FHST Updates: Scheduled : 6 Successful : 6 2.3 Significant HSTARs: None 2.4 Operations Note: The Servicing Mission Ground Test #2 (SMGT2) was successfully conducted Thursday, Jan. 28 from 8 A.M. through 1 P.M. All test objective were met. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 2257 (Physical Condition in Gasious Galactic Clouds) 3.2 Complete WFC/PC 4109 (WF/PC Calibration: Rapid Internal Monitor) 3.3 Uplink Updated Ephemeris Tables, Per MOSES 2871 3.4 Complete GHRS/2 3957 (Interstellar Molecules in Spectra of 28 Stars) 3.5 Complete PC 4347 (WF/PC Calibration: Flood Retention Test) 3.6 Complete PC 3603 (Snapshot of Protoplanetary Nebulae) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-